OpenStudy (anonymous):

find the limit as x approaches 0 for e^x cos2x

OpenStudy (anonymous):

Hiya! Start by checking two things: what does e^x become as x approaches 0, and what does cos2x become as x approaches 0?

OpenStudy (anonymous):

e^x = 0 as it approaches 0? & i don't understand how to cos2x as it approaches 0... are you supposed to substitute 0 for x as well?

OpenStudy (aum):

\[ \large \lim_{x \rightarrow 0} e^x\cos(2x) = e^0\cos(0) = ? \]